# Leveraging Photometry for Deconfusion of Directly Imaged Multi-Planet Systems

### v1

@@ 3.2 Validating deconfuser Orbit Fits
We generate the input files required by orbitize! by converting the epochs of observation to Modified Julian Day (MJD) format (JD-2400000.5). The chosen date in MJD is arbitrary, but the spacing of the observation epochs remains consistent with the epoch spacing simulated in this work. For example, MJD 59215.0, 59397.5, and 59580.0 represent three equally spaced observations spread over one year.
We initialize the MCMC sampler with 20 temperatures, 1,000 walkers, and 50,000 steps per walker. For MCMC, temperature scales the target function to adjust the efficiency of exploring the parameter space. Higher temperatures allow for more efficient exploration of the parameter space without becoming stuck in regions with high probability density; lower temperatures allow chains to efficiently sample the high probability regions (Earl & Deem, 2005; Vousden et al., 2016). We visually inspect the walkers to assess convergence and reject the first 1,000 steps before computing the posterior distribution. We use the default priors in orbitize! (see Table 1 in Blunt et al. 2020), except for the prior on semimajor axis. We reduce the maximum range on the semimajor axis prior from 10,000 AU to 10 AU to reduce computation time. This scaled-down maximum would still fall well outside the 9.7 $\lambda/D$ outer working angle (OWA) of the Roman Hybrid-Lyot Coronagraph (Riggs et al., 2021) for a system at 10 pc. The semimajor axis prior then becomes LogUniform(1,10). Of the known exoplanet systems at the time of writing, there are only 124 planets with semimajor axes greater than 10 AU, and only 79 of those have semimajor axes less than 100 AU (Christiansen et al., 2025).
Figure 7 shows the posterior distributions of orbital parameters in medium inclination system seven returned by orbitize! for planets one, two, and three, respectively. The simulated system values (“truth”) and the deconfuser fit values are shown in green and magenta, respectively, on top of the posterior histograms. The comparison between the deconfuser fits and truth values emphasize that the deconfuser is fitting the shape parameters and inclination of orbits well, with the exception of the eccentricity of the second planet. For all three planets in medium inclination system seven, the semimajor axis is fit by the deconfuser to within 0.1-26% of the true value, the eccentricity is fit to within 4-180% of the true value, and the inclination is fit within 0.9-64% of truth. The reduced accuracy of the fit of the second planet’s eccentricity by the deconfuser could be attributed to the reduced fraction of the orbit that is covered (less than half) by the detections, compared to planets one and three. In comparison, the mean of the orbitize! distributions are up to an 85% difference for the planet’s semimajor axes, up to 6700% from their true eccentricities, and up to 46% from their true inclinations. A comparison of the true orbital parameters, top ranked deconfuser fit, and orbitize! means are shown in Table 5.
Table 5: Comparison between orbitize! and deconfuser orbital parameters and their true values for medium inclination system seven. Corner plots are shown in Figure 7 and orbit solutions are shown in Figure 8. Semimajor axes ($a$) are given in units of AU, inclinations ($i$) are given in units of degrees. Deconfuser orbits were fit with a maximum tolerance of 0.05 AU.
Planet Orbit Deconfuser Truth orbitize!
no. parameter value top-ranked fit fit mean
[…]
In most cases, the fit deconfuser value for the semimajor axis, eccentricity, and inclination falls within or near the peak of the posterior distribution fit by orbitize!, validating the deconfuser’s fits to these parameters. The cases where the deconfuser fit parameter falls outside of the peak of the posterior distribution are also cases where the true parameter is outside of the distribution, and therefore orbitize! also struggled to find the true fit. For all parameters (semimajor axis, eccentricity, and inclination), the deconfuser fit falls within three sigma of the orbitize! mean. However, the orbitize! distributions are relatively widespread (especially for $a$ and $e$). The orbit fits for each of the three planets are shown in Figure 8 with the top ranked deconfuser fit (from photometry) plotted for reference.
Figure 7: Posterior distributions of orbital parameters and orbit fits estimated by MCMC with orbitize! for planets (a) one, (b) two, and (c) three in system seven. Posterior distributions are shown with the true parameters of the simulated system plotted in green and the top ranked deconfuser fit chosen by the photometry ranking in magenta. Mean values estimated by orbitize! are also listed above each histogram and in Table 5.
Figure 8: A subset of 50 random orbits fit by orbitize! (purple) for the three planets in system seven: (a) planet one, (b) planet two, and (c) planet three. The top ranked deconfuser fit returned by the new photometry ranking is shown for comparison in green. Planet detections are indicated by red stars.
4 Discussion
Here we present a more detailed discussion of the results, including how well the orbits are fit compared to the truth and the implications of the results for direct imaging missions. The results suggest that including photometry benefits deconfusion for the example systems in this work, across all inclination ranges. Across the low, medium, and high inclination cases simulated here, the photometry likelihood ranking scheme enabled the separation of confused orbit options for each of the systems. Photometry aided in choosing the better orbit fit out of the confused options for all but 30% of the cases in the low inclination group, and all but 40% of the cases in both the medium and high inclination groups.
4.1 Photometry for Deconfusion
There are several contributing factors which influence the utility of photometry for deconfusing orbit options in these example cases. The first is due to the statistical nature of photon detections when observing exoplanets and the limited number of photons which will be received when imaging at high contrast levels. The likelihoods that are used as the ranking metric in this work are derived from a distribution of potential electron counts at the detector for each planet detection, which is a combination of a Poisson distribution, Gamma distribution, and Gaussian distribution (see detailed description in Section 2.2). The combination of statistical distributions and low number of photons per observation may result in variations of estimated count rate per detection, which influences the likelihood value. This explains the potential for variation in likelihoods on a per-orbit level. That is, if there are two system options for which one planet was fit with the same orbit in both, the options will return different likelihood values between the two confused systems. This may ultimately affect the resulting system ranking for better or worse, even though all likelihood results fall within the distribution of possible expected values.
Another commonality between systems for which the “worst” option is chosen appears when planet detections are close to each other in the plane of the sky. In system one, the detections of the two inner planets appear to be less than 0.1 AU from one another in the plane of the sky at each inclination. Systems four and five also fall into this category. In system four, the second detections of the inner two planets appear closer than 0.3 AU (or 0.03′′) of one another. For this reason, the difference in likelihoods between the low inclination cases are relatively small as the predicted phase angles are similar (in combination with the limited number of received photons) and spatially close together. Figure 6 shows the likelihoods for all systems and ranked options, and is shaded based on the number of planets per ranking for which the semimajor axis was the “best” fit for the system (i.e., minimum percent difference between true and fit semimajor axes). The difficulty in separating options with detections with small separations can be seen in Figure 6 for systems one, four, and five, where the differences between system likelihoods are small. Section 5 expands on efforts that could investigate reducing the confusion caused by closely separated detections.
Figure 6 introduces some key points regarding photometry for orbit discrimination. For the example systems presented here, it is clear that photometry shows promise for supporting orbit differentiation better than geometric orbit-fitting alone. Each of the system options presented were previously closely ranked with only relative astrometric information. Likelihoods assigned to these systems by the photometry ranking scheme allow for a refined ranking of orbit combinations from the deconfuser. The photometry ranking scheme generally provides more distinction between orbits for the high inclination systems in this work, compared to lower inclination, because there is more phase variation in these observations. This can be seen by the higher scatter in likelihood values in the high inclination cases for the systems presented here (e.g., systems four, five, and nine).
4.2 The Role of Geometric Albedo
We briefly comment here on the role of geometric albedo for modeling photometry, as this is currently a constant parameter in our simulations. The choice in geometric albedo ($A_{g}$) for modeling the photometry will affect the overall likelihood values if the albedo is guessed incorrectly. A model albedo that is larger than the observed albedo would require a lower phase angle, decreased planet-star separation, or larger planet radius to produce the observed signal; a model albedo that is smaller than the observed albedo would require the opposite. An incorrectly guessed albedo might preference a fit with larger/smaller planet-star separation if the model albedo is less/greater than the truth, if such an orbit-fit has been returned by the deconfuser. However, if all other parameters are held constant, the general trend of a planet’s phase variation remains the same for all values of $A_{g}$. Because the photometry ranking scheme is choosing between and comparing orbits which have already been fit to detections, rather than influencing the orbit fitting process, we do not expect an incorrect guess for the geometric albedo to drastically change the performance of the photometry ranking.
4.3 Phase Function and Albedo Assumptions
A limitation of this work lies in the choice of a Lambertian phase function and the constant albedo to model planetary reflectance. The assumption of Lambertian reflectance approximates the shape of a planetary phase function relatively well for the type of planets and temperature regimes that would be probed by a reflected light direct imaging mission. Strong asymmetries in the reflected light curves of cool, mature planets are less likely because they are receiving relatively uniform irradiation from their host star (i.e., they are not tidally locked like known short-period transiting planets; e.g., Zellem et al. 2014; Parmentier et al. 2021). However, Solar System phase curves have been shown to deviate from Lambertian assumptions (e.g., Mayorga et al. 2016; Mallama & Hilton 2018; Mayorga et al. 2020).
Observations of Solar System planets have also demonstrated that real planetary albedos vary with wavelength and time, driven by factors such as clouds, aerosol scattering, surface features, and seasons. Mallama & Hilton (2018) discuss some of these results in detail. For example, analysis of models and observations show that the Earth may exhibit magnitude variation ranging from 0.12 to 0.76 between cloud-free and cloudy cases (e.g., Tinetti et al. 2006; Stephens et al. 2015; Mallama & Hilton 2018). The phase curve of Venus shows a reversal in brightness at high phase angles, which Mallama et al. (2006) attribute to forward scattering of sulfuric acid droplets in the atmosphere. The giant planets in the Solar System exhibit albedo variation on a smaller scale at visible wavelengths, compared to the terrestrial planets (Mallama & Hilton, 2018). Jupiter observations near full phase show magnitude variations at the level of a few hundredths due to changes in the cloud bands (Mallama & Schmude, 2012).
These complexities are not yet included in this work, but we describe future directions in Section 5 to incorporate more realistic reflectance behavior. In the scheme of the likelihood ranking, the true solution will always lie in the sample of possible solutions that are calculated, if the forward model assumptions are correct. However, one might reject the best-fit orbit solution if the model lacks fidelity. For example, in the case where the true planetary reflectance is non-Lambertian.
4.4 Exo-moons and Exo-rings
Other factors that may influence exoplanetary photometry are moons and rings, which could have a negative impact on the photometry likelihood ranking scheme. Rings alter the spectrum of a planet and would introduce asymmetries to the phase curve in reflected light (Arnold & Schneider, 2004; Dyudina et al., 2005; Coulter et al., 2022). Arnold & Schneider (2004) noted that an icy ring around an Earth-like planet could result in a brightness 100 times that of a ringless planet. In such a case, additional observations at longer wavelengths would be important for breaking the albedo-radius degeneracy and inferring the presence of rings.
For exomoons, Limbach et al. (2024) showed that the influence on the spectrum of an Earth-like planet is minimal near visible wavelengths. Their work concluded that detecting a moon-Earth mutual event at 550 nm with a bandwidth of 20% with an 8-meter class telescope would require around two to twenty mutual events. Such cases where observations measure anomalous photometry due to some influencing factors would need to be considered on a case-by-case basis. Using a more complex phase function and incorporating additional filters or polarimetry would likely be more beneficial for ranking such systems where phase curve asymmetries are being induced by material in the circumplanetary environment.
4.5 Effect of Planetary Motion on Astrometry and Photometry
The motion of a planet around its star and through the obscured regions of an observing instrument will affect the ability to measure a planet’s position and photometry. This depends to some extent on the orbit parameters of the planet and the observing system. If a planet goes inside the IWA or outside the OWA (outer-working angle) of the observing instrument, it will not be possible to get accurate astrometry or measure photometry. If the planet has an eccentric orbit around its star, the cloud and surface properties may change substantially with orbital position. For example, varying instellation will cause different types of clouds to condense at different regions along the planet’s orbit (e.g., Marley et al. 2013; Cahoy et al. 2010). It may also cause the surface to freeze over or thaw, therefore increasing or decreasing the surface albedo if the orbit takes a planet in/out of cooler and warmer conditions (e.g., Venkatesan et al. 2025).
P22 also show that missed detections (e.g., when a planet falls inside/outside the working angles of the instrument) are likely to increase confusion, because this may increase the number of orbits that can be fit to the points that are visible to the observer. However, where a planet’s photometry can be measured, the likelihood ranking scheme would operate in the same manner. For the current model of photometric behavior and radiometric brightness, these keep-out zones and portions of the orbit where planets are not observable would not be input to the deconfuser. The user would only provide the observable points, and thus, the deconfuser would only fit orbits to the input points (P22). This would not change any of the outcomes of this work, as a user would have measured photometry for the input points that are provided to the deconfuser. Section 5 identifies future work to address cases where planets are unobservable in their orbit.
4.6 Comparing Photometry Ranking to Dynamical Assumptions
A significant fraction of known exoplanet systems exhibit eccentricities greater than 0.1. Of the systems with measured eccentricities, 46% have $e>0.1$ and 27% have $e>0.2$ (Christiansen et al., 2025). Therefore, while dynamical stability is an important and necessary criterion for deconfusion, photometry provides a critical, independent discriminator for deconfusion decisions.
4.7 Observing Strategy
As was presented in P22, the level of confusion depends on the number of observations and the system architecture. With three visits, the probability of confusion is still relatively high for high inclination systems ($40\%$; P22). Increasing the number of visits to a system to four or five can significantly reduce ambiguity, but this comes at a cost to overall mission yield (e.g., Stark et al. 2016; Morgan et al. 2022). Each observation of a system requires hours of integration time (dependent on the host star’s brightness and the required contrast) that use up mission resources and time that could be spent characterizing other systems. Due to these trade-offs, it is beneficial to overall mission yield to reduce the number of observations needed to characterize a system/planet and optimize the observing strategy for each system. Efficient deconfusion that leverages all available information from a series of observations is highly valuable in this regard.
4.8 Accuracy of Deconfuser Orbit Fits
To perform a thorough assessment of the deconfuser’s performance, we also calculated the accuracy of the semimajor axis fits returned by the deconfuser compared to the true semimajor axis for each planet, by comparing their percent difference. The accuracy of the fit for the semimajor axis is important for enabling follow-up observations and characterization of planets. The tolerance for orbital parameter accuracy defined by the NASA Standards Definition and Evaluation Team for Yield Analysis (Morgan et al., 2019) is to better than ten percent accuracy. Of the planets in the low inclination systems simulated here, 64 percent of the top ranked orbit fits retrieved a semimajor axis within five percent of the true semimajor axis, 77 percent were retrieved to within ten percent of the true semimajor axis. All planets except one were fit to an accuracy of better than 25 percent. Of the medium inclination systems, 40 percent of the orbit fits retrieved the semimajor axis to within five percent of the true value, 70 percent to within ten percent of the true value, and 93 percent to within 25 percent of the true value. For the high inclination systems, the number of planets for which semimajor axes were retrieved to better than five percent of the true value dropped to 37 percent, 57 percent for constraints within ten percent, and 73 percent for constraints within 25 percent of the true semimajor axis. We emphasize that these percentages result after selection of the “best” orbit using the new photometry ranking scheme, where the best orbit is the one with the minimum percent difference between individual orbital parameters compared to the true system. In order to characterize potentially habitable planets with a future mission like HWO, observation schedulers will require accurate knowledge of whether a planet lies within the host star’s potential habitable zone and where it will be in the future in order to schedule the next observation.
4.9 Importance of Deconfusion for Observation Planning
Being able to deconfuse multiple orbit options using photometry will ultimately improve observation scheduling and enable characterization of exoplanet systems. It will also inform decisions about how to allocate mission time to studying planets of interest. For example, the confused orbit options for the inner two planets in the high inclination system ten give fits that change the number of planets with separations that fall within the boundaries of a Solar-type star’s liquid-water habitable zone (defined here with an inner edge at 0.836 AU and outer edge at 1.656 AU as in Kane & Gelino 2012). In the first option for the high inclination system ten, only the second planet in the system would fall within the liquid-water habitable zone. In the second option, both planet two and planet three would fall within the habitable zone. The difference between these orbital separations and architectures may inform how we determine follow-up observations on such a system. Observers may choose not to allocate valuable mission observing time and resources in a scenario where an orbit option places a planet outside of the habitable zone. The deconfuser with its photometry ranking scheme can support this decision making.
5 Summary & Future Work
In this work, we introduce a method for using orbital phase and planetary brightness variation to deconfuse orbit combinations. We present a straightforward analysis of a limited set of cases to demonstrate the utility of planetary phase for planet-detection deconfusion. We apply the updated orbit ranking scheme to thirty systems containing three planets, categorized into three groups of ten across low, medium, and high inclination groupings. We adopt some instrument parameters for the Roman Space Telescope Coronagraph Instrument to implement the photometric orbit-ranking scheme with realistic instrument values.
The results presented here show that combining photometry with astrometry shows promise for improving deconfusion of multi-planet systems and enabling orbit differentiation. For the example systems presented here, each had multiple closely ranked orbit options using only relative astrometry, which results in confusion. The photometry ranking scheme helps to choose the system fit that most closely resembles the true simulated system in all but 30% of the ten low inclination cases, all but 40% of the ten medium inclination cases, and all but 40% of the ten high inclination cases presented here. Future work will explore the potential for improving the error rates on orbit ranking with photometry by setting thresholds for comparing system likelihoods and using more realistic planetary reflectance treatments.
On top of supporting observation scheduling and follow-up decisions by providing an estimate of where a planet resides in its system, orbit determination also supports placing better constraints on exoplanet properties. Salvador et al. (2024) showed that a priori knowledge of a planet’s orbit supports accurate retrieval of the radius in reflected light observations. Orbit determination helps to break the degeneracy between planet radius and phase angle. Carrión-González et al. (2020) concluded that knowing a planet’s radius can help differentiate between clear and cloudy atmospheric scenarios, and supports placing constraints on atmospheric abundances of species like CH4. An unknown planet radius reduces the accuracy of the atmospheric retrieval for reflected light spectroscopy. Therefore, reducing the radius uncertainty on a planet through orbit determination is a useful prior for exoplanet characterization.
We also validated the orbits fit by the deconfuser by presenting the first detailed comparison to MCMC fits from the orbitize! software. Results indicate that the true parameters and top deconfuser fits after ranking with photometry for the semimajor axis, eccentricity, and inclination parameters typically fall within the posterior distributions returned by orbitize!. The deconfuser fit parameters for semimajor axis, eccentricity, and inclination are compared to the orbitize! mean values and the truth values in Table 5. Instances where the deconfuser’s fits fall outside of the peak of the posterior distribution are also those where the true parameter falls outside of the peak, indicating that orbitize! also struggled to find the correct fit.
During direct imaging observations, planets may also be blocked by the inner or outer working angles of the observing instrument. This would result in missed detections when the planet falls within these obscured regions. Detections might also be missed if they are at a phase angle that is too dim to observe. In these cases, the probability of confusion for a system will be higher (P22). However, with photometric information, we may be able to reject or assign a lower ranking to geometrically valid orbits which would require the planet to be at a phase angle which is inconsistent with the detection’s brightness.
Future work will also account for the joint dependence of astrometry and photometry in the simulations. That is, the knowledge of a planet’s astrometry will be less accurate if its phase angle is higher and the planet is therefore dimmer, and vice versa. A detailed study of the effect of non-detections on confusion rates and the scheduling (and number) of visits per system would also directly inform yield calculations for future missions. Adjusting the observing cadence of a system may support reducing confusion that remains due to closely separated planet detections. Deviating from uniformly spaced observations, which were previously suggested for single-planet systems (Horning et al., 2019), might enable better distinction in separation and phase space. With additional orbit ranking metrics and simulations, we can provide a more complete assessment of confusion rates and their impact on yield for a mission like HWO.
Acknowledgements
S. N. H. would like to thank Eric Nielsen for helpful discussion on MCMC orbit-fitting with orbitize! and Vanessa Bailey for support in specifying the Roman Coronagraph Instrument parameters. The authors would like to thank the reviewers for their helpful feedback and suggestions. This research was carried out at the Massachusetts Institute of Technology and in part at the Jet Propulsion Laboratory, California Institute of Technology.
Appendix A Likelihood Ranking
Here we describe the likelihood ranking scheme developed for deconfusion with photometry. In this work, we consider each planet detection to be an independent event, as described in Section 2.2. Each planet detection will have a unique photon count that is dependent on factors including the orbital parameters of the system, the observation time, the background contribution, and the planet’s properties. The number of photons detected also depends on the parameters of the observing instrument, which motivates the inclusion of the detector model described in Sections 2.1 and 2.2.
Here we describe the number of counts collected by the observing detector with a slightly simplified approach for clarity, adopting the observation time as $\Delta t$, the background counts as $d$, and the photon flux as $\phi$. The background here includes contributions from both the detector and astrophysical sources. The expected number of photons for a detection ($n_{p}$) is obtained by combining these possible contributions to the photon count, which can be described by Equation A1.
$n_{p}(a,e,i,\Omega,\omega,M_{0})=\phi_{p}(a,e,i,\Omega,\omega,M_{0})\cdot\Delta t+d$ (A1)
The orbital parameter definitions are listed in Table 4.
@@ Appendix A Likelihood Ranking
$L_{\mathrm{detection}}=f(y|n_{p}(a,e,i,\Omega,\omega,M_{0})),$ (A2)
where $f(y|n_{p})$ is the distribution returned by the noise model as described in Section 2.2. This distribution represents the measured photons from the combination of the planet detection and detector noise. Each set of planet detections must be considered together as part of a system in order to determine the likelihood of the whole system’s orbit matches ($L_{\mathrm{system}}$). The likelihood of measuring all detections for a system given the parameters of the matched orbits can be described by the product of the individual detections (Equation A3),
$L_{\mathrm{system}}=\prod_{j=1}^{N}f(y_{j}|n_{p,j}(a_{j},e_{j},i_{j},\Omega_{j},\omega_{j},M_{0,j})),$ (A3)
@@ Appendix B The Solar System as a Confused System
Table 6: Orbital parameters of the simulated Solar System planets (Venus, Earth, Mars, Jupiter) and the number of confused orbit options per system. Detections and confused orbits are shown in Figures 9 and 10 for the medium and high inclination cases, respectively.
@@
We also apply the photometry likelihood ranking scheme to a group of Solar System planets as if they were a directly-imaged exoplanetary system. We have chosen to model the four planets that fall within the semimajor axis range defined for the rest of the simulated systems in this paper, which includes Venus, Earth, Mars, and Jupiter. Orbital parameters were adopted from JPL Horizons Solar System Dynamics page [^5] https://ssd.jpl.nasa.gov/planets/approx_pos.html (Standish et al., 1992). We follow the same simulation methods as are described in Section 2.3, with three observations of the system spread equally over the span of one Earth year. Venus, Earth, Mars, and Jupiter do not exhibit confusion at low inclinations. We present a medium and high inclination case here.
We maintain the range of the true inclinations of the planets with respect to one another and increase the overall viewing inclination by randomly sampling an inclination value in the medium and high inclination groups and incorporating it for all planets. The orbital parameters are listed in Table 6. We assign albedos and radii in the same manner as presented in Section 2.3.2 above, assuming albedos of 0.3 and Earth-like radii. We also tested the photometry model using the true geometric albedos and radii of the planets; the results are the same as the case with the constant Earth-like albedos and radii. Changing the albedos and radii of planets in this Solar System case does not change the orbits that are fit by the deconfuser (because photometry does not influence the orbit-fitting portion of the algorithm). It also does not affect the results of photometry ranking, because the majority of the confusion in the Solar System cases stems from the detections of Earth and Venus being at small separations relative to one another on-sky.
In both the medium and high inclination scenarios, the deconfuser returns two confused orbit options for the four Solar System planets. The medium inclination detections and confused options are shown in Figure 9. The high inclination detections and confused options are shown in Figure 10.
The top-ranked orbit options are compared to the true orbits in Figure 11. The confused orbit options and their rankings are given in Table 7. In the medium inclination case, the photometry ranking scheme ranks option A (panel b in Figure 9) higher than option B (panel c in Figure 9). Option A also contains the orbits which have the best match to the true system. That is, the semimajor axes, eccentricities, and inclinations in option A have the smallest percent difference between the two options, compared to the true system.
In the high inclination case, the photometry ranking scheme ranks option B (panel c in Figure 10) as the top option, despite it being the “worst” fit of the two options. The statistical nature of photon detections compound with the small separations between Venus and Earth in this scenario, as discussed in Section 4. The likelihood of the orbit fit for Jupiter also varies between options one and two, despite the parameters being the same, due to the limited number of photons received.
The results of this Solar System test agree with the results from the simulated systems presented in Sections 3.1 and 4.1 above. Photometry largely shows promise for deconfusion of otherwise closely ranked orbit options, but can be complicated by low-photon statistics and small separations between multiple planets in a system.
@@
Table 7: Solar system photometry ranking results. Orbital parameters and updated rankings for the confused orbit options returned by the deconfuser for the medium, and high inclination Solar system case. The confused orbit options are shown in Figures 9 and 10.
“Confused” orbit parameters
Incl. System Planet $a$ $e$ $i$ Deconfuser RMS Orbit System System

### v2

@@ Appendix A Likelihood Ranking
$L_{\mathrm{detection}}=f(y|n_{p}(a,e,i,\Omega,\omega,M_{0})),$ (10)
where $f(y|n_{p})$ is the distribution returned by the noise model as described in Sec. 2.2. This distribution represents the measured photons from the combination of the planet detection and detector noise. Each set of planet detections must be considered together as part of a system in order to determine the likelihood of the whole system’s orbit matches ($L_{\mathrm{system}}$). The likelihood of measuring all detections for a system given the parameters of the matched orbits can be described by the product of the individual detections,
$L_{\mathrm{system}}=\prod_{j=1}^{N}f(y_{j}|n_{p,j}(a_{j},e_{j},i_{j},\Omega_{j},\omega_{j},M_{0,j})),$ (11)
@@ Appendix B The Solar System as a Confused System
We also apply the photometry likelihood ranking scheme to a group of Solar System planets as if they were a directly-imaged exoplanetary system. We have chosen to model the four planets that fall within the semimajor axis range defined for the rest of the simulated systems in this paper, which includes Venus, Earth, Mars, and Jupiter. Orbital parameters were adopted from JPL Horizons Solar System Dynamics page (https://ssd.jpl.nasa.gov/planets/approx_pos.html) [61]. We follow the same simulation methods as are described in Sec. 2.3, with three observations of the system spread equally over the span of one Earth year. Venus, Earth, Mars, and Jupiter do not exhibit confusion at low inclinations. We present a medium and high inclination case here.
We maintain the range of the true inclinations of the planets with respect to one another and increase the overall viewing inclination by randomly sampling an inclination value in the medium and high inclination groups and incorporating it for all planets. The orbital parameters are listed in Table 5. We assign albedos and radii in the same manner as presented in Sec. 2.3.2 above, assuming albedos of 0.3 and Earth-like radii. We also tested the photometry model using the true geometric albedos and radii of the planets; the results are the same as the case with the constant Earth-like albedos and radii. Changing the albedos and radii of planets in this Solar System case does not change the orbits that are fit by the deconfuser (because photometry does not influence the orbit-fitting portion of the algorithm). It also does not affect the results of photometry ranking, because the majority of the confusion in the Solar System cases stems from the detections of Earth and Venus being at small separations relative to one another on-sky.
Table 5: Orbital parameters of the simulated Solar System planets (Venus, Earth, Mars, Jupiter) and the number of confused orbit options per system. Detections and confused orbits are shown in Figs. 7 and 8 for the medium and high inclination cases, respectively.
@@
In both the medium and high inclination scenarios, the deconfuser returns two confused orbit options for the four Solar System planets. The medium inclination detections and confused options are shown in Fig. 7. The high inclination detections and confused options are shown in Fig. 8.
@@
The top-ranked orbit options are compared to the true orbits in Fig. 9. The confused orbit options and their rankings are given in Table 6. In the medium inclination case, the photometry ranking scheme ranks option A (panel b in Fig. 7) higher than option B (panel c in Fig. 7). Option A also contains the orbits which have the best match to the true system. That is, the semimajor axes, eccentricities, and inclinations in option A have the smallest percent difference between the two options, compared to the true system.
(a)                   (b)
@@
Table 6: Solar system photometry ranking results. Orbital parameters and updated rankings for the confused orbit options returned by the deconfuser for the medium, and high inclination Solar system case. The confused orbit options are shown in Figs. 7 and 8.
“Confused” orbit parameters Deconfuser
Incl. System Planet no. $a$ $e$ $i$ RMS Orbit System System
@@
In the high inclination case, the photometry ranking scheme ranks option B (panel c in Fig. 8) as the top option, despite it being the “worst” fit of the two options. The statistical nature of photon detections compound with the small separations between Venus and Earth in this scenario, as discussed in Sec. 4. The likelihood of the orbit fit for Jupiter also varies between options one and two, despite the parameters being the same, due to the limited number of photons received.
The results of this Solar System test agree with the results from the simulated systems presented in Secs. 3.1 and 4.1 above. Photometry largely shows promise for deconfusion of otherwise closely ranked orbit options, but can be complicated by low-photon statistics and small separations between multiple planets in a system.
Disclosures
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper.
Code, Data, and Materials Availability
The code developed in this manuscript is available on GitHub: https://github.com/MIT-STARLab/deconfuser. Orbital parameters for the systems described in this work are provided in supplemental machine-readable files.